# Beyond Black-Box Interventions: Latent Probing for Faithful Retrieval-Augmented Generation

### 13 Apr 2026 v2

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often fail to maintain contextual faithfulness, generating responses that conflict with the provided context or fail to fully leverage the provided evidence. Existing methods attempt to improve faithfulness through external interventions, such as specialized prompting, decoding-based calibration, or preference optimization. However, since these approaches treat the LLM as a black box, they lack a reliable mechanism to assess when and why knowledge conflicts occur. Consequently, they tend to be brittle, data-intensive, and agnostic to the model’s internal reasoning process. In this paper, we move beyond black-box interventions to analyze the model’s internal reasoning process. We discover that conflicting and aligned knowledge states are linearly separable in the model’s latent space, and contextual noise systematically increases the entropy of these representations. Based on these findings, we propose ProbeRAG, a novel framework for faithful RAG that operates in three stages: (i) fine-grained knowledge pruning to filter irrelevant context, (ii) latent conflict probing to identify hard conflicts in the model’s latent space, and (iii) conflict-aware attention to modulate attention heads toward faithful context integration. Extensive experiments demonstrate that ProbeRAG substantially improves both accuracy and contextual faithfulness. The related resources are available at https://github.com/LinfengGao/ProbeRAG.
@@
Linfeng Gao1,4thanks:    Work done during internship at Alibaba Cloud Intelligence.   Qinggang Zhang2   Baolong Bi3   Bo Zeng4   Zheng Yuan2   Zerui Chen1 Zhimin Wei1Shenghua Liu3Linlong Xu4Longyue Wang4Weihua Luo4Jinsong Su1thanks:    Corresponding author 1School of Informatics, Xiamen University 2The Hong Kong Polytechnic University 3University of Chinese Academy of Sciences 4Alibaba Cloud Intelligence gaolinfeng@stu.xmu.edu.cn, qinggang.zhang@connect.polyu.hk, jssu@xmu.edu.cn
@@ 5 Related Work
Retrieval-Augmented Generation (RAG) has emerged as a prominent paradigm for enhancing the factual accuracy and temporal relevance of Large Language Models (LLMs) by incorporating external knowledge sources (Shi et al., 2024; Xiang et al., 2026; Chen et al., 2025b; Xiao et al., 2025; Hui et al., 2025; Chen et al., 2025c). Early works such as REALM (Guu et al., 2020a) and RAG (Lewis et al., 2020) retrieve relevant passages from large corpora to assist generation. Subsequent research has explored improvements in both the retriever and generator modules, including dense retrieval techniques (Karpukhin et al., 2020; Izacard et al., 2023), adaptive retrieval strategies (Sun et al., 2022; Chen et al., 2025a), and hybrid models combining retrieval with parametric memory (Shi et al., 2023b; Wang et al., 2025e).
@@ Appendix F Related Work
Contextual faithfulness, defined as the degree to which model outputs remain consistent with retrieved or provided context (Niu et al., 2024; Wang et al., 2025c), has emerged as a central concern in RAG research (Malaviya et al., 2023; Chern et al., 2023; Li et al., 2024). Without explicit mechanisms to enforce faithfulness, models may hallucinate, overgeneralize, or generate outputs inconsistent with retrieved passages (Chen et al., 2024b; Sui et al., 2025).
@@
Beyond methodological innovations, recent surveys (Zhou et al., 2023b; Ji et al., 2023) highlight persistent challenges in faithfulness evaluation. Automatic metrics such as factual consistency (Thorne et al., 2018) or entailment-based scores (Falke et al., 2019; Guo et al., 2023) provide useful proxies but often fail to capture nuanced inconsistencies or omissions. Consequently, many works advocate for human-in-the-loop evaluation frameworks to assess contextual grounding at scale.
@@
Recent research also extends conflict resolution beyond the text domain. Multimodal RAG systems (Gao et al., 2023; Xu et al., 2024b) face analogous challenges, as retrieved visual or audio evidence may not align with textual outputs. This motivates broader frameworks for consistency checking across modalities. Furthermore, evaluation efforts (Xu et al., 2024a) emphasize the need for standardized benchmarks that explicitly include conflict scenarios, enabling more systematic analysis of models’ conflict-handling behaviors.
@@ Appendix H Reproducibility Statement
Our code, datasets, and implementation details are anonymously available at https://github.com/LinfengGao/ProbeRAG. We make significant efforts to ensure the reproducibility of our work. The details of model architectures, hyperparameters, and training settings are provided in Section 4.1 of the main paper. Additional implementation details and full experimental setups are provided in Appendix C.

### 15 Apr 2026 v3

@@ Abstract
Retrieval-Augmented Generation (RAG) systems often fail to maintain contextual faithfulness, generating responses that conflict with the provided context or fail to fully leverage the provided evidence. Existing methods attempt to improve faithfulness through external interventions, such as specialized prompting, decoding-based calibration, or preference optimization. However, since these approaches treat the LLM as a black box, they lack a reliable mechanism to assess when and why knowledge conflicts occur. Consequently, they tend to be brittle, data-intensive, and agnostic to the model’s internal reasoning process. In this paper, we move beyond black-box interventions to analyze the model’s internal reasoning process. We discover that conflicting and aligned knowledge states are linearly separable in the model’s latent space, and contextual noise systematically increases the entropy of these representations. Based on these findings, we propose ProbeRAG, a novel framework for faithful RAG that operates in three stages: (i) fine-grained knowledge pruning to filter irrelevant context, (ii) latent conflict probing to identify hard conflicts in the model’s latent space, and (iii) conflict-aware attention to modulate attention heads toward faithful context integration. Extensive experiments demonstrate that ProbeRAG substantially improves both accuracy and contextual faithfulness. The related resources are available at https://github.com/XMUDeepLIT/ProbeRAG.
@@
Linfeng Gao1,4thanks:    Work done during internship at Alibaba Group.   Qinggang Zhang2thanks:    Corresponding author   Baolong Bi3   Bo Zeng4   Zheng Yuan2   Zerui Chen1 Zhimin Wei1Shenghua Liu3Linlong Xu4Longyue Wang4Weihua Luo4Jinsong Su1,522footnotemark: 2 1School of Informatics, Xiamen University 2The Hong Kong Polytechnic University 3University of Chinese Academy of Sciences 4Alibaba Group 5Key Laboratory of Digital Protection and Intelligent Processing of Intangible Cultural Heritage of Fujian and Taiwan (Xiamen University), Ministry of Culture and Tourism, China gaolinfeng@stu.xmu.edu.cn, qinggang.zhang@connect.polyu.hk, jssu@xmu.edu.cn
@@ 5 Related Work
Retrieval-Augmented Generation (RAG) has emerged as a prominent paradigm for enhancing the factual accuracy and temporal relevance of Large Language Models (LLMs) by incorporating external knowledge sources (Shi et al., 2024; Xiang et al., 2026a; Chen et al., 2025b; Xiao et al., 2025; Hui et al., 2025; Chen et al., 2025c). Even before the emergence of LLMs, researchers were already using RAG technology to improve various natural language processing tasks, such as machine translation (Jiang et al., 2022; Cao et al., 2023; Gao et al., 2024). Early works such as REALM (Guu et al., 2020a) and RAG (Lewis et al., 2020) retrieve relevant passages from large corpora to assist generation. Subsequent research has explored improvements in both the retriever and generator modules, including dense retrieval techniques (Karpukhin et al., 2020; Izacard et al., 2023), adaptive retrieval strategies (Sun et al., 2022; Chen et al., 2025a), and hybrid models combining retrieval with parametric memory (Shi et al., 2023b; Cao et al., 2024; Wang et al., 2025e).
@@ 6 Conclusion
Acknowledgements
The project was supported by National Key R&D Program of China (No. 2022ZD0160501), Natural Science Foundation of Fujian Province of China (No. 2024J011001), and the Public Technology Service Platform Project of Xiamen (No.3502Z20231043). We also thank the reviewers for their insightful comments.
@@ Appendix F Related Work
Contextual faithfulness, defined as the degree to which model outputs remain consistent with retrieved or provided context (Niu et al., 2024; Wang et al., 2025c), has emerged as a central concern in RAG research (Chern et al., 2023; Li et al., 2024). Without explicit mechanisms to enforce faithfulness, models may hallucinate, overgeneralize, or generate outputs inconsistent with retrieved passages (Chen et al., 2024b; Sui et al., 2025).
@@
Beyond methodological innovations, recent surveys (Zhou et al., 2023b; Ji et al., 2023; Xiang et al., 2026b) highlight persistent challenges in faithfulness evaluation. Automatic metrics such as factual consistency (Thorne et al., 2018) or entailment-based scores (Falke et al., 2019; Guo et al., 2023) provide useful proxies but often fail to capture nuanced inconsistencies or omissions. Consequently, many works advocate for human-in-the-loop evaluation frameworks to assess contextual grounding at scale.
@@
Recent research also extends conflict resolution beyond the text domain. Multimodal RAG systems (Gao et al., 2023; Xu et al., 2024b; Yang et al., 2026) face analogous challenges, as retrieved visual or audio evidence may not align with textual outputs. This motivates broader frameworks for consistency checking across modalities. Furthermore, evaluation efforts (Xu et al., 2024a) emphasize the need for standardized benchmarks that explicitly include conflict scenarios, enabling more systematic analysis of models’ conflict-handling behaviors.
@@ Appendix H Reproducibility Statement
Our code, datasets, and implementation details are anonymously available at https://github.com/XMUDeepLIT/ProbeRAG. We make significant efforts to ensure the reproducibility of our work. The details of model architectures, hyperparameters, and training settings are provided in Section 4.1 of the main paper. Additional implementation details and full experimental setups are provided in Appendix C.